OpenStudy (anonymous):

for the purpose of purchasing new baseboard and carpet. a. find the area and perimeter of the room (neglecting doors). b.indentify whether baseboard has to do with area or perimeter and the same with carpet. the lenght is10.5ft and the weight is 9ft. a. the area of the room is _________sq ft and the perimeter is _________-ft.?

OpenStudy (stormfire1):

I assume you mean the length is 10.5ft and the width is 9ft. A = L * W so the area is 10.5 * 9 = 94.5ft^2 P = 2L + 2W = 2(10.5) + 2(9) = 21 + 18 = 39ft The baseboard has to do with the perimeter since baseboards go around the room.

OpenStudy (anonymous):

does the carpet has to do with the area? nd thankk u for showing your work

OpenStudy (stormfire1):

Yes, since the carpet must cover the entire floor you need to know the area of the room.

OpenStudy (anonymous):

thankk u.

OpenStudy (stormfire1):

np :)